OpenStudy (hungergames13):

Which ordered pair is the solution to the system of equations? -3x+4y=-20 x+4y=-4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (gabbyalicorn):

This is what I got: x+4y=−4;−3x+4y=−20 Step: Solvex+4y=−4for x: x+4y+−4y=−4+−4y(Add -4y to both sides) x=−4y−4 Step: Substitute−4y−4forxin−3x+4y=−20: −3x+4y=−20 −3(−4y−4)+4y=−20 16y+12=−20(Simplify both sides of the equation) 16y+12+−12=−20+−12(Add -12 to both sides) 16y=−32 16y16=−3216(Divide both sides by 16) y=−2 Step: Substitute−2foryinx=−4y−4: x=−4y−4 x=(−4)(−2)−4 x=4(Simplify both sides of the equation) Answer: x=4 and y=−2